OpenStudy (anonymous):

find the inverse function of f(x)= 5/x-2

OpenStudy (anonymous):

f inverse = (5/x) + 2

OpenStudy (anonymous):

how you did it?

OpenStudy (anonymous):

type y = 5/ x-2 then make x the subject of the formula

OpenStudy (anonymous):

where did you get Y=5, and the other. Can you do it step by step plz?

OpenStudy (anonymous):

is it \[f(x)=\frac{5}{x}-2\] or \[f(x)=\frac{5}{x-2}\] ?

OpenStudy (anonymous):

steps are 1) y = 5/(x-2) 2) y(x-2) = 5 3) x= (5/y)+2 4) change x to y and y to x making it y = (5/x) +2

OpenStudy (anonymous):

the last one you wrote

OpenStudy (anonymous):

i am confuse , between the second to the third step

OpenStudy (anonymous):

from second to third step you (x-2) = 5/y then x=(5/y) + 2